OpenStudy (anonymous):

You are going to write a fairy tale! Your story should include all of the standard elements of fairy tales. Begins “Once upon a time…” Animals can act like humans Magical items or helpers Ends “…happily ever after.” Your story must also fully cover these math concepts. This is an educational fairy tale… the best kind! Create a quadratic function with two real zeros. Solve a system of non-linear equations with a graph. Solve a system of non-linear equations with a table. Compare linear, quadratic, and exponential growth functions using a table and a graph.

OpenStudy (anonymous):

can someone PLEASE please help. I'm trying to finish my Florida virtual and only need this done and one more assignment. I'm lost on it.

OpenStudy (anonymous):

Once upon a time a student with and English question wandered into the mathematics section of openstudy.com....

OpenStudy (usukidoll):

ROFL

OpenStudy (anonymous):

Thanks, math fans. Make a quadratic real zeroes a, b by multiplying (x-a)(x-b)=0

OpenStudy (anonymous):

Solve a system of non-linear equations with a graph f = sin x g = cos x g - f = 0

OpenStudy (anonymous):

y = mx + b linear y = ax^2 + c quadratic y = exp(x) exponential